# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

### v2

@@ Abstract
On-device skin lesion edge classification remains hindered by the lack of efficient architectures compatible with neuromorphic deployment, as conventional CNN-to-SNN conversion suffers from structural incompatibility, performance degradation on rare diseases, and absence of continual learning. To address these limitations, we propose an end-to-end pipeline that incorporates a quantization-aware CNN backbone explicitly designed for conversion-preserving mapping to SNNs and neuromorphic deployment, culminating in the QANA architecture optimized for low-latency, energy-efficient, and incremental inference. QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference. Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms. Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison. Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods. These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
I Introduction
Skin diseases present significant diagnostic challenges, particularly for clinicians without specialized training, leading to frequent misdiagnosis of conditions such as melanoma and Merkel cell carcinoma [1, 2]. Although deep learning-based diagnostic systems have shown promising performance [3], most approaches rely on centralized training and inference, requiring sensitive patient data to be transferred to cloud servers [4]. This raises data security risks [5] and is constrained by strict privacy regulations such as HIPAA [6] and GDPR [7]. Furthermore, to support dermatological care in home and remote settings lacking conventional healthcare infrastructure[8], it is essential to develop models that enable on-device training and inference.
However, deploying deep learning models on edge devices faces several challenges. Conventional CNNs require significant computational and energy resources, increasing device complexity and limiting portability [9]. In addition, effective training typically depends on large labeled datasets, which are often unavailable for rare skin diseases [10, 11]. Frequent model updates further introduce computational overhead and privacy concerns, while hardware constraints restrict model complexity and inference accuracy on edge platforms [2].
@@ I Introduction
Figure 1: Detailed architecture of our end-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a novel quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
Recently, Spiking Neural Networks (SNNs) and neuromorphic computing platforms have emerged as promising alternatives for overcoming deployment challenges in dermatological diagnosis [12]. Unlike conventional CNNs, SNNs use discrete spike events for information encoding and transmission, resulting in sparse and event-driven computation that greatly reduces power consumption [13]. This spike-based encoding is well-suited for scenarios with limited labeled data, as the temporal dynamics of spikes facilitate effective learning from fewer examples [14]. Neuromorphic processors, such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17], natively support SNNs and enable on-chip incremental learning. Incremental learning enables these systems to adapt efficiently to new patient data without full model retraining, aligning well with clinical practices that regularly acquire new diagnostic cases [18]. Additionally, these neuromorphic devices possess compact physical footprints and substantially lower energy requirements compared to conventional GPUs [19], making them viable candidates for edge-based deployment in portable diagnostic instruments.
The prevalent approach for neuromorphic deployment is to first train conventional CNN architectures, such as ResNet or DenseNet, on large datasets, and then convert them to equivalent SNN models for edge inference [20]. While this conversion aims to combine the accuracy of CNNs with the efficiency of SNNs [21], several practical limitations persist. Key CNN components, such as Batch Norm, Flatten, and Global Average Pooling, cannot be directly mapped to spike-based neuron units, complicating and often degrading the conversion process [22]. In addition, converted SNNs frequently experience significant accuracy loss on small, imbalanced medical datasets due to quantization effects and limited capacity for capturing subtle pathological features [23].
We propose a quantization-aware neuromorphic architecture (QANA) within an end-to-end pipeline (Fig. 1) to overcome these challenge. QANA comprises three essential modules: (1) iterative feature extraction and downsampling block utilizing stacked Ghost modules, spatially-aware efficient channel attention (SA-ECA), and residual connections to reduce computation and inference latency, thereby directly facilitating effective CNN-to-SNN conversion; (2) spike-compatible feature transformation block employing quantization-aware normalization, bounded activations, and squeeze-and-excitation (SE) blocks, ensuring hardware compatibility and efficient incremental learning; and (3) quantized output projection block explicitly optimized for conversion-preserving CNN-to-SNN mapping and low-power neuromorphic deployment. These components collectively address key conversion limitations, preserve high accuracy for rare diseases, and efficiently support incremental model updates on edge devices.
Evaluations on the HAM10000 benchmark dataset show that QANA achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%. Further validation on a clinical dataset confirms robustness, achieving a Top-1 accuracy of 90.8% and a macro F1 score of 81.7%. When deployed on the Akida neuromorphic hardware platform, QANA attains an inference latency of 1.5 ms per image and energy consumption of 1.7 mJ, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, surpassing state-of-the-art CNN-to-SNN conversion methods in computational efficiency. These outcomes verify QANA’s practical suitability for resource-constrained and privacy-sensitive medical diagnostic applications.
Our main contributions are:
  • Conversion-preserving architecture. We design QANA with bounded activations, separable and ghost convolutions, and lightweight channel attention with SE gating, ensuring all operators lower one-to-one to 8-bit spiking kernels.
  • Deployment-aligned training and conversion. We provide a graph-preserving flow with batch-norm fusion and affine 8-bit quantization, combined with QAT, embedding-space SMOTE with safety filtering, and a lightweight on-chip incremental readout update, all under a fixed $64{\times}64$ input and matched preprocessing.
  • Edge evaluation. On Akida and under identical preprocessing and quantization for baselines, QANA achieves 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical set, with 1.5 ms latency and 1.7 mJ per image.
Figure 2: Detailed architecture of our QANA, which performs iterative feature extraction using stacked Ghost modules, ECA, and residual blocks, followed by spike-compatible transformation with batch normalization, ReLU activation, and Squeeze-and-Excitation (SE) block. The output is then quantized and projected to class logits for SNN deployment.
II Related Work
This section surveys the landscape for edge dermatology. We first outline neuromorphic hardware capabilities and constraints that govern feasible model design. We then compare ANN-to-SNN conversion with direct SNN training, including spike encoding and sparsity control, and finally review CNN-based dermatology systems as accuracy baselines. Across these threads we emphasize limitations under small and imbalanced datasets and strict latency and energy budgets, which together define the problem setting.
II-A Neuromorphic Hardware for Edge Imaging
Neuromorphic processors execute event-driven computation with low power and support on-chip adaptation, which matches the requirements of on-device medical imaging [14]. Akida, TrueNorth, and Loihi provide hardware support for spiking inference and incremental learning [15, 16, 17]. At the same time, these platforms constrain model design through limited operator sets, fixed-point quantization, and resource budgets that restrict input resolution and intermediate state. Related methods must therefore account for quantization and operator compatibility during both training and deployment.
II-B ANN-to-SNN Conversion
A prevalent strategy trains a conventional CNN and converts it to an SNN by mapping activations to spikes with calibrated thresholds and quantized ranges [24]. Conversion enables reuse of mature backbones but introduces several sources of error. Batch normalization and global pooling do not admit direct spike-based equivalents, which forces structural edits and alters feature statistics. Rate-coded readouts require temporal integration that increases latency and makes predictions sensitive to threshold and timestep calibration, which is difficult to tune on small and imbalanced datasets. Empirical studies report measurable accuracy drops after conversion under these conditions, indicating that quantization and operator mismatch are primary failure modes on edge workloads [15].
II-C Directly Trained SNNs and Spike Encoding
Direct training replaces the non-differentiable spike with surrogate derivatives and optimizes spiking dynamics end-to-end. SLAYER derives a time-domain error assignment rule for convolutional SNNs with learnable delays [25], and surrogate-gradient frameworks formalize stable optimization for deep networks [26]. Subsequent work improves few-timestep accuracy with residual spiking blocks [27] and temporal efficient training that couples the loss across discrete integration steps [28]. Despite these advances, training remains sensitive to surrogate choice, thresholds, and timestep schedules. Performance depends on spike encoding. Time-to-first-spike and related temporal codes reduce spikes and latency [29, 30], yet they complicate calibration on hardware designed for rate accumulation and may reduce robustness on fine-grained classes when data are scarce. Activity regularization controls event rates and energy [31, 25], but aggressive sparsity can underfit rare patterns. These factors make it difficult to guarantee accuracy, latency, and energy simultaneously under fixed hardware constraints.
II-D CNN-Based Dermatology Models
Dermatology classifiers based on DenseNet-121 and Inception-v4 reach strong results on ISIC-style datasets [32, 33]. In practice, deployment often depends on cloud retraining to handle distribution shift and rare categories, which introduces latency and privacy risk [18, 34]. Training data are class-imbalanced and images vary across devices and skin tones, which limits generalization to underrepresented conditions [35]. These limitations indicate that comparative evaluation on edge hardware should consider both conversion-based pipelines and directly trained SNNs with temporal encodings and sparsity control, so that the impact of quantization, operator compatibility, and encoding choices on accuracy, latency, and energy can be isolated under identical deployment settings.
III Methodology
@@ III-A Data Preprocessing
III-A1 Image Preprocessing
All dermatoscopic images undergo automated integrity checks followed by a brief manual review when needed. Inputs are processed at $64\times 64$ to match the Akida on-chip memory budget used in our deployment. To retain lesion content at this resolution, we apply Grad-CAM on the pretrained backbone to obtain a coarse saliency map. A tight square crop is centered on the top-activated region and resized with antialiased area resampling to $64\times 64$. If the saliency peak lies within 10% of the image center or the saliency map is flat (coefficient of variation <0.05<0.05), we fall back to a centered square crop. Pixel intensities are normalized per channel to $[0,1]$. No denoising or inpainting is used.
III-A2 Data Augmentation
@@
III-A3 Synthetic Minority Oversampling (SMOTE)
@@
where $\boldsymbol{\mu}$ is the training mean, $\mathbf{U}_{d}\in\mathbb{R}^{4096\times d}$ the leading eigenvectors and $\mathbf{\Lambda}_{d}\in\mathbb{R}^{d\times d}$ their eigenvalues. Nearest neighbors are computed with Euclidean distance in the whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
@@
$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ (2)
$\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$
@@
Synthetic data are not images. They are injected at the penultimate layer as additional feature vectors and used to train the quantized output projection with cross-entropy. During this stage the backbone up to the spike-compatible transform is frozen; only the classifier head is updated for 5 epochs with mixed real and synthetic embeddings. A subsequent short fine-tuning on real data only re-enables the last block and the head to reconcile the feature distribution. Input images are still resized to $64\times 64$ prior to feature extraction for hardware constraints, but no pixel-space mixing or image synthesis is performed at any time.
III-B Quantization-Aware Network Architecture
The model (as shown in Fig. 2) comprises a hierarchical cascade of Ghost-based multi-scale feature extraction blocks, integrated channel attention mechanisms, and a quantization-aware transformation stage that produces SNN-ready outputs. This architecture balances computational efficiency and discriminative power, facilitating direct conversion and robust inference on the Akida neuromorphic platform.
III-B1 Iterative Feature Extraction and Downsampling
Feature extraction begins with a stack of Ghost blocks [36], designed to maximize representational diversity while minimizing computational cost. As shown in Fig. 3, each Ghost block receives a feature tensor $F^{(l-1)}$ (for $l=1,\ldots,4$) and applies both base and ghost convolutions as:
$F_{\text{ghost}}^{(l)}=\text{Concat}\Big($ $\underbrace{\mathcal{F}_{1\times 1}^{(l)}\big(F^{(l-1)}\big)}_{\text{base: pointwise conv}},\,\underbrace{\mathcal{F}_{3\times 3}^{(l)}\big(F^{(l-1)}\big)\odot\mathbf{M}^{(l)}}_{\text{ghost: depthwise conv with filter mask}}\Big)$ (6)
Here, $\mathcal{F}_{k\times k}^{(l)}$ is a separable convolution of kernel size $k\times k$, and $\mathbf{M}^{(l)}$ is a binary mask selecting learnable ghost filters. This combination generates both local and extended receptive fields, enabling richer features with fewer FLOPs.
After concatenation, the tensor is processed by batch normalization, quantization-bounded activation, and dropout:
$F_{\text{drop}}^{(l)}=\text{Dropout}\Big(\min\big(6,\,\max(0,\,\gamma^{(l)}\cdot\text{BN}(F_{\text{ghost}}^{(l)})+\beta^{(l)})\big)\Big)$ (7)
This quantization-aware step is crucial for downstream SNN compatibility and robust generalization.
To capture cross-channel correlations, each block integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) mechanism, inspired by Efficient Channel Attention (ECA) mechanism [37]. As shown in Fig. 4, Instead of computing global channel statistics via global average pooling and 1D convolution, we adopt a lightweight depthwise convolution followed by pointwise channel-wise scaling. This enables efficient modeling of spatial-channel dependencies with minimal overhead and neuromorphic compatibility:
$\widetilde{F}^{(l)}=\sigma\left(\mathbf{W}_{1\times 1}^{(l)}*\text{BN}\left(\text{DWConv}_{k\times k}\left(F^{(l)}\right)\right)\right)\odot F^{(l)}$ (8)
where $\text{DWConv}_{k\times k}$ and $\mathbf{W}_{1\times 1}^{(l)}$ jointly form the attention mechanism in our SA-ECA block, and $\sigma$ is the sigmoid activation used to generate the attention mask. This lightweight attention preserves channel expressiveness with minimal overhead.
To ensure stable deep stacking, we employ residual skip connections and spatial downsampling:
$F^{(l)}=\text{MaxPool2D}\left(\boldsymbol{\alpha}^{(l)}\odot F_{\text{drop}}^{(l)}+\mathbf{P}^{(l)}F^{(l-1)}\right)$ (9)
where $\mathbf{P}^{(l)}$ projects the previous block’s output for dimension alignment if needed. This design preserves gradient flow, reduces vanishing/exploding risk, and aggregates multi-scale context, all critical for reliable feature extraction in small-data regimes. After four such blocks, the spatial size is reduced to $4\times 4$.
@@ III-B1 Iterative Feature Extraction and Downsampling
Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block. A depthwise convolution is first applied to extract channel-wise statistics, followed by a lightweight 1D convolution to model local channel dependencies. The resulting attention weights are used to rescale the input feature channels, enhancing discriminative information with minimal computational overhead.
III-B2 Spike-Compatible Feature Transformation
The output of the previous stage, $F^{(4)}$, is passed to a spike-compatible transformation module engineered for direct quantization and SNN integration. A SeparableConv2D $(3\times 3,256)$ [38] generates higher-dimensional features, followed by quantization-aware normalization and bounded activation:
$\widehat{F}=\min\left(6,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (10)
This mapping guarantees that all activations lie in $[0,1]$, which is both compatible with spike encoding and preserves information for subsequent inference.
To further optimize the channel-wise information flow, as shown in Fig. 5, a Squeeze-and-Excitation (SE) block [39] is applied, implementing a two-stage bottleneck and gating mechanism:
$\mathbf{s}=\sigma\left(W_{2}\,\delta\left(W_{1}\frac{1}{16}\sum_{i,j}\widehat{F}(i,j,:)\right)\right)$ (11)
where $W_{1}$ and $W_{2}$ are dense layers, $\delta$ is ReLU, and $\sigma$ is sigmoid. Each feature channel is then scaled by $s_{c}$, which adaptively modulates discriminative capacity and provides additional regularization for small, imbalanced datasets.

### v3

# Quantization-Aware Neuromorphic Architecture for Skin Disease Classification on Resource-Constrained Devices

@@ Abstract
On-device skin lesion analysis is constrained by the compute and energy cost of conventional CNN inference and by the need to update models as new patient data become available. Neuromorphic processors provide event-driven sparse computation and support on-chip incremental learning, yet deployment is often hindered by CNN-to-SNN conversion failures, including non-spike-compatible operators and accuracy degradation under class imbalance. We propose QANA, a quantization-aware CNN backbone embedded in an end-to-end pipeline engineered for conversion-stable neuromorphic execution. QANA replaces conversion-fragile components with spike-compatible transformations by bounding intermediate activations and aligning normalization with low-bit quantization, reducing conversion-induced distortion that disproportionately impacts rare classes. Efficiency is achieved through Ghost-based feature generation under tight FLOP budgets, while spatially-aware efficient channel attention and squeeze-and-excitation recalibrate channels without heavy global operators that are difficult to map to spiking cores. The resulting quantized projection head produces SNN-ready logits and enables incremental updates on edge hardware without full retraining or data offloading. On HAM10000, QANA achieves 91.6% Top-1 accuracy and 91.0% macro F1, improving the strongest converted SNN baseline by 3.5 percentage points in Top-1 accuracy, corresponding to a 4.0% relative gain, and by 12.0 points in macro F1, corresponding to a 15.2% relative gain. On a clinical dataset, QANA achieves 90.8% Top-1 accuracy and 81.7% macro F1, improving the strongest converted SNN baseline by 3.2 points in Top-1 accuracy, which corresponds to a 3.7% relative gain, and by 3.6 points in macro F1, corresponding to a 4.6% relative gain. When deployed on BrainChip Akida, QANA runs in 1.5 ms per image with 1.7 mJ per image, corresponding to 94.6% lower latency and 99.0% lower energy than its GPU-based CNN implementation. Source code is available at Akida-Research-Project.
1 Introduction
Skin lesions include benign conditions and malignant disease, and timely recognition is essential for early intervention and appropriate referral. In Australia, melanoma is among the most frequently diagnosed cancers, with 10,600 cases in males and 7,600 cases in females in 2023 [2]. Cancer Council Australia further notes that Australia has one of the highest melanoma rates worldwide [6]. This high burden increases the number of lesions requiring assessment and amplifies the clinical cost of delayed detection and incorrect triage. Because melanoma can closely resemble benign lesions in dermatoscopic appearance, there is a practical need for portable decision-support systems that perform multi-class triage at the point of care and help distinguish melanoma from visually similar benign conditions. Although deep learning-based diagnostic systems have achieved strong performance [3], many deployments remain cloud-centric and require transferring sensitive patient data [43], which increases security risk [21] and is constrained by privacy regulation, including HIPAA [24] and GDPR [22]. For home and remote settings where connectivity and infrastructure are limited [16], practical systems must support on-device inference and efficient model adaptation without repeated cloud retraining.
@@ 1 Introduction
Figure 1: End-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
Recently, Spiking Neural Networks (SNNs) and neuromorphic hardware have been explored for edge dermatological diagnosis [29]. SNNs encode information as spike events, enabling sparse and event-driven computation with low power consumption [10]. Their temporal coding can be effective in limited-label regimes [18]. Neuromorphic processors such as BrainChip Akida [5], IBM TrueNorth [4], and Intel Loihi [34] support SNN execution and on-chip incremental learning, allowing updates as new cases arrive [7]. Their compact form factor and reduced energy demand relative to GPUs further support portable, privacy-preserving deployment [14]. A common route to neuromorphic deployment is to train a conventional CNN backbone such as ResNet or DenseNet and convert it to an SNN for edge inference [33, 27]. In practice, conversion is often brittle because standard CNN components such as BatchNorm and global pooling are not naturally expressed in spike-based neuron operators, leading to operator substitution and calibration sensitivity [19]. Accuracy can further degrade on small, imbalanced medical datasets due to quantization effects and reduced capacity to preserve rare lesion cues [13].
We propose a Quantization Aware Neuromorphic Architecture (QANA) within the end-to-end pipeline as depicted in Fig. 1 to address these challenges. QANA consists of three modules. First, an iterative feature extraction and downsampling stage uses stacked Ghost blocks, spatially-aware efficient channel attention, and residual connections to reduce compute while preserving discriminative features required for dermatology. Second, a spike-compatible transformation stage applies quantization-aware normalization, bounded activations, and squeeze-and-excitation to produce activations suitable for spike encoding and hardware constraints. Third, a quantized projection head maps the compact representation to class logits in a form that remains compatible with CNN-to-SNN conversion and low-power neuromorphic execution.
On HAM10000, QANA achieves 91.6% Top-1 accuracy and 91.0% macro F1, improving the best converted SNN baseline by 3.5 points in Top-1 accuracy and 12.0 points in macro F1. On the clinical dataset, QANA attains 90.8% Top-1 accuracy and 81.7% macro F1, improving over the strongest baseline by 3.2 points in Top-1 accuracy and 3.6 points in macro F1. On BrainChip Akida, QANA processes an image in 1.5 ms at 1.7 mJ, yielding 94.6% lower latency and 99.0% lower energy than its CNN implementation on an RTX 3090.
The primary contributions of this research are as follows:
  • We design QANA, a conversion-oriented quantization-aware architecture for dermatoscopic image classification that replaces conversion-fragile components with spike-compatible, bounded-activation blocks and lightweight attention to preserve discriminative features under 8-bit constraints.
  • We formulate a graph-preserving CNN-to-SNN conversion pipeline for Akida that integrates BatchNorm reparameterization, uniform affine integer quantization, and deterministic operator lowering to ensure an all-integer executable spiking graph. Source code is available at Akida-Research-Project.
  • We develop a deployment and adaptation procedure on neuromorphic hardware that calibrates spike thresholds and integration windows and supports incremental updates by fine-tuning only the readout layer to accommodate domain shift without full retraining.
2 Related Work
Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [18]. Platforms such as BrainChip’s Akida [5], IBM’s TrueNorth [4], and Intel’s Loihi [34] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining. These features are advantageous for medical imaging tasks on edge devices. However, existing works primarily focus on general benchmarks like MNIST [26] or CIFAR [42] and rarely address domain-specific challenges in dermatology [36]. Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [17].
CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures, including ResNet and DenseNet, for deployment on neuromorphic processors [28]. Conversion toolkits, including Akida’s CNN2SNN [39], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents. While some studies report preserved accuracy on standard datasets [35], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [5]. Operations such as batch normalization and multi-bit activations require careful folding and scale management in integer-only conversion pipelines, and they are a frequent source of calibration error that can amplify performance loss on imbalanced medical data [41].
In parallel, conventional CNN-based methods, such as those using DenseNet-121 [8] or Inception-v4 [37], show high performance on benchmark datasets like ISIC. Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [7]. Such retraining incurs latency, high computational cost, and data privacy concerns [25]. Moreover, general CNN models often fail to generalize to underrepresented conditions [30], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions. These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
3 Methodology
@@ 3.1 Data Preprocessing
3.1.1 Image Preprocessing
All dermatoscopic images undergo automated integrity checks followed by a brief manual review. Inputs are processed at $64\times 64$ to match the Akida on-chip memory budget used in our deployment. To retain lesion content at this resolution, we apply Grad-CAM on the pretrained backbone to obtain a coarse saliency map. A tight square crop is centered on the top-activated region and resized with antialiased area resampling to $64\times 64$. If the saliency peak lies within 10% of the image center or the saliency map is flat (coefficient of variation <0.05<0.05), we fall back to a centered square crop. Pixel intensities are normalized per channel to $[0,1]$. No denoising or inpainting is used. Cropping is computed per image without using class labels. The saliency model used to produce the Grad-CAM map is fixed before training, and the same cropping rule is applied to training, validation, and test images. This ensures that the preprocessing is deterministic given an input image and does not introduce split-dependent adaptation. For Grad-CAM generation we use a fixed ResNet-50 pretrained on ImageNet. This network is not fine-tuned on HAM10000 or the clinical dataset and is used only to produce a saliency prior for cropping. Since cropping depends only on the input image and a frozen saliency model, it does not use labels and it does not incorporate any information from validation or test outcomes.
3.1.2 Data Augmentation
@@
Figure 2: Architecture of QANA. The backbone performs iterative feature extraction using stacked Ghost modules [11], spatially-aware efficient channel attention [40], and residual blocks. A spike-compatible transformation stage applies bounded activation and squeeze-and-excitation [12] before quantized projection to class logits.
3.1.3 Synthetic Minority Oversampling (SMOTE)
@@
where $\boldsymbol{\mu}$ is the training mean, $\mathbf{U}_{d}\in\mathbb{R}^{4096\times d}$ the leading eigenvectors and $\mathbf{\Lambda}_{d}\in\mathbb{R}^{d\times d}$ their eigenvalues. Nearest neighbors are computed with Euclidean distance in whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
@@
$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ $\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$ (2)
@@
Synthetic data are not images. They are injected at the penultimate layer as additional feature vectors and used to train the quantized output projection with cross-entropy. During this stage the backbone up to the spike-compatible transform is frozen; only the classifier head is updated for 5 epochs with mixed real and synthetic embeddings. A subsequent short fine-tuning on real data re-enables the last block and the head to reconcile the feature distribution. Input images are still resized to $64\times 64$ prior to feature extraction for hardware constraints, but no pixel-space mixing or image synthesis is performed at any time.
@@
Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block. A depthwise convolution is applied to capture lightweight spatial-channel context, followed by batch normalization and a $1\times 1$ pointwise projection to produce attention weights. The resulting weights rescale the input feature channels with minimal computational overhead.
3.2 Quantization-Aware Network Architecture
As shown in Fig. 2, QANA follows a hierarchical design with four lightweight feature extraction stages and a spike-compatible transformation stage that produces SNN-ready activations for deterministic CNN-to-SNN conversion on Akida. The backbone prioritizes conversion stability by constraining activation ranges and by using operators that admit integer lowering, while preserving discriminative capacity through multi-scale feature generation and channel recalibration.
Backbone configuration. The input is $64\times 64\times 3$. The backbone contains four stages, each consisting of a Ghost block followed by BN, ReLU6, and a residual shortcut. Spatial downsampling is performed by $2\times 2$ max pooling with stride $2$ at the end of each stage, yielding a $4\times 4$ feature map after stage four. When the shortcut changes the channel dimension, a $1\times 1$ projection is used to match dimensions before residual addition. Dropout is applied only during training after the bounded activation to reduce overfitting on minority classes.
3.2.1 Iterative Feature Extraction and Downsampling
Feature extraction is implemented as a stack of Ghost blocks [11] that expand representational diversity under a tight FLOP budget. For stage $l\in\{1,\ldots,4\}$, the input tensor $F^{(l-1)}$ is first mapped to a compact set of primary features using a pointwise projection. Additional features are then generated using depthwise transformations, and the two parts are concatenated to form the output feature map:
$F_{\mathrm{p}}^{(l)}$ $=\mathcal{F}_{1\times 1}^{(l)}\!\left(F^{(l-1)}\right),F_{\mathrm{g}}^{(l)}$ $=\mathcal{F}_{\mathrm{dw}}^{(l)}\!\left(F_{\mathrm{p}}^{(l)}\right),F_{\mathrm{ghost}}^{(l)}$ $=\mathrm{Concat}\!\left(F_{\mathrm{p}}^{(l)},\,F_{\mathrm{g}}^{(l)}\right).$ (6)
In practice, the channel allocation between $F_{\mathrm{p}}^{(l)}$ and $F_{\mathrm{g}}^{(l)}$ follows the Ghost design ratio used in Fig. 3, ensuring that most computation is concentrated in the primary projection while the remaining channels are produced with depthwise operations. After feature generation, each stage applies BatchNorm followed by a bounded ReLU6 activation to keep post-activation values within a fixed and quantization-friendly range, then applies dropout for regularization. This bounded design reduces calibration sensitivity during 8-bit quantization and improves conversion robustness when activations are later encoded into spikes.
To model channel dependencies without introducing conversion-fragile global operators, each stage integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) block inspired by ECA [40]. As illustrated in Fig. 4, SA-ECA replaces global pooling and 1D channel convolution with a depthwise spatial aggregation and a $1\times 1$ projection that produces per-channel rescaling weights, followed by a sigmoid gate. The resulting weights modulate the stage features through channel-wise multiplication, yielding a lightweight attention mechanism that is compatible with integer operator lowering.
To maintain stable deep stacking in small-data regimes, each stage uses a residual shortcut with optional projection for dimension alignment, followed by spatial downsampling via max pooling. After four stages, the spatial resolution is reduced to $4\times 4$, producing a compact feature tensor suitable for spike-compatible transformation.